# Flow of Reasoning: Training LLMs for Divergent Problem Solving with Minimal Examples

### v4

@@ 4.6.1 Results
4.7 Mathematical Reasoning: GSM8K
Table 6: Results on GSM8K.
Method Acc. (%) Diversity
CoT (2-shot) 45.72 1.12
CoT-SC (2-shot) 41.74 -
RAP 37.16 -
SFT ($\alpha=1.0$) 52.69 1.13
FoR 57.39 1.26
Setup. We use the last 50 training examples in the original training dataset, and we sample 4 times for every problem at inference. We follow the baseline implementations from LLM reasoners (Hao et al., 2024). Since there is no standard metric for reasoning diversity, we manually annotate 50 test examples to evaluate semantic difference between solutions as the diversity metric.
Reward Design. We removed the success reward since the pretrained mathematical knowledge enables LLMs to perform relatively well compared to BlocksWorld. Using LLMs to verify each state is computationally expensive. The reward is defined as $R(s_{n})=w\cdot\mathbb{I}(\text{success})$.
@@ 4.7 Mathematical Reasoning: GSM8K
FoR (Ours) 100.00 (0.00) 98.41 (1.12) 1.27 (0.02) 78.44 (4.54) 1.33 (0.03)
4.8 Additional Analysis
Ablation Study. To further demonstrate the effectiveness of FoR, we conduct ablation studies to analyze the impact of individual components, focusing specifically on the BlocksWorld task. Table 7 summarizes the experimental results. 1) "Local search" significantly enhances the performance of FoR by improving exploration in the trajectory space and collecting high-reward trajectories for training. Removing local search results in a 31.3% decrease in 6-step task accuracy. 2) $\epsilon$-sampling also contributes to exploration, though to a lesser extent. 3) "Augmented intermediate rewards" play a critical role, as removing them leads to a 51% drop in 6-step accuracy and a 13% reduction in diversity. The left plot in Figure 3 shows the impact of varying the augmented reward weight $\lambda$. This improvement arises from distinguishing unsuccessful trajectories that are more likely and less likely to succeed, guiding the policy towards more successful paths by assigning higher probabilities to the former. However, when $\lambda$ becomes too large, accuracy declines as the success reward’s influence diminishes. 4) the replay buffer contributes to FoR performance by leveraging historical high-reward trajectories for learning.
Figure 3: Additional Analysis on BlocksWorld. Left: Accuracy of FoR across different step settings with varying intermediate reward weight ($\lambda$). Right: Comparison of accuracy and diversity between SFT trained with varying data sizes and FoR trained with 15 examples on the test dataset.
Data-Efficiency. We tested the final 20 examples from the 6-step test set and adjusted the number of training examples for both FoR and the SFT methods. As illustrated in the right plot of Figure 3, SFT’s accuracy improves with additional training data, while its diversity remains stable. However, SFT’s performance remains lower than FoR’s under any amount of training data. This is attributed to FoR’s ability to learn from divergent reasoning trajectories, enhancing trajectory coverage and improving generalization to new cases.
Discussion. FoR achieves higher reasoning accuracy and diversity by leveraging its exploratory nature (madan2024goalflownet) to sample multiple valid trajectories, covering a broader solution space without getting stuck on incorrect paths like SFT. This increases the likelihood of reaching a correct solution. A detailed case study that compares FoR with SFT is provided in Appendix F.
@@ 5 Conclusion
We introduce FoR that efficiently trains LLM policy for diverse, high-quality reasoning paths with probability proportional to unnormalized reward. The core of the approach is the flow-based formulation of multi-step reasoning that allows us to adapt principled GFlowNet training strategies. On five representative tasks across embodied, math, logical, spatial, and abstraction reasoning, FoR show stronger performance and improved diversity than both finetuning-based and prompting-based baselines. We discuss limitations and broader impact in Appendix G and H.
Acknowledgement
We would like to thank Jalend Bantupalli, Yanze Lu, and Yitong Gu for their assistance with the coding aspects of this work.
@@ Appendix A Additional Related Work
Recent LLMs (Achiam et al., 2023; Touvron et al., 2023; Bai et al., 2022; Chowdhery et al., 2023) have demonstrated great potentials in tackling complex reasoning tasks (Cobbe et al., 2021; Mishra et al., 2022; Hendrycks et al., 2021; Rein et al., 2023; Mialon et al., 2023). (1) Fine-tuning LLMs is a primary way to enhance their reasoning abilities, including SFT and reinforcement learning (RL) approaches. SFT with large-scale and high-quality datasets of reasoning chains has proven very effective (Yu et al., 2023c; Yue et al., 2023; Yuan et al., 2024a). Various methods for constructing training samples have been proposed when ground truth reasoning chains are not available. For example, STaR (Zelikman et al., 2022) uses online sampling with self-correction to find positive samples. ReSTEM (Singh et al., 2023) and V-STaR(Hosseini et al., 2024) filter samples with external verifiers. RL techniques, particularly reward-maximizing policy optimization methods like PPO, are widely employed in LLMs (Ouyang et al., 2022; Bai et al., 2022; Havrilla et al., 2024; Luong et al., 2024). However, both maximum likelihood training (i.e. SFT) and reward-maximizing policy optimization (e.g., PPO) do not encourage models to generate diverse solutions. (2) prompting-based reasoning algorithms aim to better elicit the knowledge inside LLMs without tuning their parameters. Techniques such as CoT (Wei et al., 2022) and its variants (Chen et al., 2022; Li et al., 2024b; Zhang et al., 2023e; Zhou et al., 2022; Kojima et al., 2022) have improved LLM performance by enabling them to generate intermediate steps before arriving at a final answer. To provide reasoning more guidance , self-evaluation (Xie et al., 2024b; Shinn et al., 2024; Madaan et al., 2024) and reward models are introduced to enhance reasoning process (Uesato et al., 2022; Lightman et al., 2023) Besides, a more relevant series of works combine LLM reasoning capabilities with planning and search algorithms such as MCTS (Hao et al., 2023; Feng et al., 2023; Zhao et al., 2024), tree and graph search (Jung et al., 2022; Zhu et al., 2022; Yao et al., 2024; Besta et al., 2024; Yao et al., 2023). Moreover, recent studies turn to amortizing computation for reasoning paths (Yuan et al., 2023; Wu et al., 2024; Bansal et al., 2024), such as self-correct (Kumar et al., 2024; Saunders et al., 2022), self-improvement (Tian et al., 2024; Yuan et al., 2024b).
GFlowNets.
@@
GFlowNets (Bengio et al., 2021) were originally proposed to learn policies for sampling from unnormalized distributions, with a primary motivation from scientific discovery (Jain et al., 2023a), which requires generating diverse high-reward samples (Shen et al., 2023b; Roy et al., 2023; Zhang et al., 2023c; Ma et al., 2024; Pan et al., 2023b), such as molecular generation (Koziarski et al., 2024; Kim et al., 2024a; Lu et al., 2024) and biological sequence generation (Ghari et al., 2023; Jain et al., 2022). Beyond the science domain, GFlowNets have also been applied in various downstream applications such as recommendation systems (Liu et al., 2023b), domain adaptation (Zhu et al., 2023), combinatorial optimization (Zhang et al., 2023b; Kim et al., 2024b) and explainability of deep neural networks (Li et al., 2023b). Additionally, GFlowNets have proven to be suitable for sampling from posterior distributions (Hu et al., 2023b; Deleu et al., 2022; 2024; Zhang et al., 2022). As a reinforcement learning method, prior works have incorporated intermediate feedback with GFlowNets to address sparse reward issues (Pan et al., 2023a; Jang et al., 2023; Pan et al., 2022) and multi-objective rewards (Jain et al., 2023b; Hernandez-Garcia et al., 2023; Chen & Mauch, 2023). There are also theoretical analyses treating GFlowNets as recurrent MCMC (Deleu & Bengio, 2023) and variational inference (Malkin et al., 2022b; Zimmermann et al., 2022) that are used to model the distribution over trajectories.
Lateral and vertical thinking.
@@
Vertical and lateral thinking (Waks, 1997; Ismayilzada et al., 2024) are two distinct approaches that differ significantly in their focus and methodology. Vertical thinking emphasizes logical, structured, and sequential reasoning, often following a step-by-step approach to solve problems. Our work aligns with this paradigm to generate multiple correct, structured reasoning processes to achieve specific goals. In contrast, lateral thinking prioritizes creativity and innovation, encouraging the exploration of unconventional perspectives and challenging established assumptions. Multiple benchmarks are proposed to evaluate the lateral thinking ability of LLMs (Huang et al., 2023; Chen et al., 2024a; Kraaijveld et al., 2024; Todd et al., 2024). To further improve the ability of lateral and divergent thinking, (Zhong et al., 2024) designs a fine-tuning and inference framework to generate unexpectable but reasonable answers, and  (Summers-Stay et al., 2023) proposes a prompting framework to enhance such ability. In addition, riddle-solving QA tasks that require reasoning about unexpected or unconventional answers such as BrainTeaser(Jiang et al., 2023) and RiddleSense (Lin et al., 2021). Future work should investigate formalizing these tasks and developing quantitative approaches to effectively guide LLMs in tackling them.
@@ Appendix B Preliminaries and Background
GFlowNets (Bengio et al., 2021; 2023; Liu et al., 2023a) are a class of models that amortize the cost of sampling from an intractable target distribution over terminal states $\mathcal{X}$ by learning a neural network-facilitated approximation of the target distribution using its unnormalized density or reward function. The task of sampling from this distribution resorts to a decision-making process. Below we introduce GFlowNets with more details.
Settings.
We are given a pointed directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$, where $S$ is a finite set of vertices (states), and $\mathcal{A}\subseteq\mathcal{S}\times\mathcal{S}$ is a set of directed edges (actions). If $s\to s^{\prime}$ is an action, we say $s$ is a parent of $s^{\prime}$ and $s^{\prime}$ is a child of $s$. There is exactly one state that has no incoming edge, called the initial state $s_{0}\in S$. States that have no outgoing edges are called terminal states. We denote by $\mathcal{X}$ the set of terminal states. A complete trajectory is a sequence $\tau=(s_{0}\to\ldots\to s_{n})$ such that each $s_{i}\to s_{i+1}$ is an action and $s_{n}=x\in\mathcal{X}$. We denote by $\mathcal{T}$ the set of complete trajectories and the terminal state as $\tau_{x}$.
@@
Training.
Training GFlowNets considers achieving a consistent flow (Bengio et al., 2023; Malkin et al., 2022a), which means the flow for the forward direction should equal to the flow for the backward direction. Below we introduce relevant objectives.
Detailed Balance (DB)
The DB objective (Bengio et al., 2023) requires learning two objectives in addition to parametric forward policy $P_{F}(\cdot|s)$: 1. A Backward policy, which is distribution $P_{B}(s^{\prime}|s;\theta)$ over the parents of any non-initial state. 2. A State flow function: $F(\cdot;\theta):\mathcal{S}\to\mathbb{R}_{>0}$. Then DB loss for a single transition $s\to s^{\prime}$ is defined as:
@@
Trajectory Balance (TB)
Trajectory balance (Malkin et al., 2022a) introduces a backward policy $P_{B}$, which is a learned distribution $P_{B}(\cdot|s^{\prime})$ over the parents of every state $s\in S$, and an estimated partition function $Z_{\theta}$ that is a scalar parameter describes the flow of initial state $F{s_{0}}$ in the DB loss. The TB objective for a complete trajectory $\tau$ is defined as
@@
If $\mathcal{L}_{TB}$ is made equal to 0 for every complete trajectory $\tau$, then 9 satisfies for all $x\in X$ and $Z$ is the inverse constant of proportionality: $Z=\sum_{x\in\mathcal{X}}R(x)$.
Conditional GFlowNets.
In a GFlowNet, both the policy and reward function can be conditioned on additional information. For instance, in the tasks we focus on, a GFlowNet policy generates actions sequentially for an embodied reasoning problem, starting from an initial state $s_{0}$ and a goal $g$. Furthermore, the allowable actions vary depending on the specific $s_{0}$ in each case. The conditional GFlowNets we train achieve amortization by sharing the policy model across different $s_{0}$ and $g$, enabling the model to generalize to initial states and targets that were not seen during training.
@@ C.3 Efficiency Analysis
FoR Setup.
During the training, we finetune the LLM with LoRA (Hu et al., 2021) with $r=32$, $\alpha=64$, and dropout=0.1. We set $\epsilon$ from 0.3 and decrease it to 0.01, $\beta$ from 1 to 2, and the probability $\delta$ using replay buffer increases from 0.3 to 0.5 throughout the iterations linearly. The learning rate is set to 1e-4 with a cosine annealing schedule, and the number of training iterations is set to 10. Reward weight $\lambda$ is set to 1.5. In our ablation study when setting $\lambda=0$, we add a small number $b=0.5$ to avoid $\log 0$. Table 9 shows the template we use for the forward policy in the 6-step setting, and its difference between 2-step and 4-step is only replacing the 6-step demonstration to 2-step and 4-step. During testing, we sample 8, 20, and 40 trajectories for 2, 4, and 6 steps respectively. As long as one trajectory reaches the goal, we label this instance as solved, all the baselines conform to the same rule.
Additional details for baselines.
We compare FoR the following baselines:
@@
(5) Proximal Policy Optimization (PPO) (Schulman et al., 2017): This is a widely-used reinforcement learning method for LLM training. We design the objective to encourage the LLM to generate solutions that satisfy the goal. Following the common practice of previous work (Ouyang et al., 2022; Wang et al., 2023a), we penalize the policy if it deviates too much from the reference policy. Formally, the objective is $\max_{\pi_{\theta}}\mathbb{E}_{\tau\sim\pi_{\theta}}\left[R(x,y)\right]-\beta\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}(y\mid x)\|\pi_{\text{ref }}(y\mid x)\right]$.
@@
CoT (1-shot) 9.52 1.0 3.12 2.02 1.0 0
CoT (5-shot) 14.28 1.0 3.12 12.12 1.08 3.45
CoT (15-shot) 11.90 1.0 3.12 8.08 1.0 0
@@
Table 10: Baseline results with diversity-encouraging instruction prompt.
@@
Performance on OOD settings.
We further assess performance on out-of-distribution (OOD) settings. Specifically, we train the model using FoR and SFT on a 2-step training set and evaluate them on a 4-step test set, and train the model on the 4-step training set and evaluate them on the 6-step test set. This allows us to analyze their generalization on OOD problems. For prompting-based baselines, we use 2-step and 4-step examples as demonstrations, respectively.
@@
Additional baseline results with diversity-encouraging instruction.
To further stimulate the diverse problem-solving ability in the baseline approaches, we add a diversity-encouraging prompt as instruction:
@@
Additional case study.
In Table 11, we show examples generated by FoR. We observe that after training, FoR can sample the terminal state with probability approximately proportional to the rewards, leading to an approximate sampling of different plans with the same probability. This empirically verifies the efficacy of the training objective.
@@
FoR Setup.
See Figure 10 for the prompt template used in the experiment of the Game of 24.
We use LoRA to train the model with $r=8$, $\alpha=32$, dropout=0.1. We load the LLM in fp16, and set the hyperparameters as follows: batch size = 4, learning rate = 1e-5, number of epochs = 5, and the reward weight $w=100$.

### v6

@@ 4.7.1 Results
4.8 Additional Analysis
Ablation Study. To further demonstrate the effectiveness of FoR, we conduct ablation studies to analyze the impact of individual components, focusing specifically on the BlocksWorld task. Table 5 and Table 8 summarize our component-wise ablations:
Table 8: Ablation results of FoR on BlocksWorld without local search.
@@ 4.8 Additional Analysis
Ours - w/o local search 100.00 89.68 1.18 53.90 1.31
- w/o augmented rewards 100.00 71.43 1.04 30.30 1.17
- w/o replay buffer 100.00 78.57 1.10 34.34 1.15
- w/o $\epsilon$-sampling 100.00 83.33 1.12 49.49 1.14
1) Local search enhances exploration and high-reward trajectory collection—its removal causes a 31.3% absolute drop in 6-step accuracy; 2) $\epsilon$-sampling also aids exploration, though to a lesser degree; 3) Augmented intermediate rewards are critical—removing them yields a 51% decrease in 6-step accuracy and a 13% reduction in diversity, and tuning the reward weight $\lambda$ (left plot of Fig. 3) further refines performance (overly large $\lambda$ harms accuracy); 4) The replay buffer leverages past high-reward trajectories—dropping it incurs a 4–6% loss with local search versus an 11–19% loss without, highlighting its greater importance when search is absent. To isolate these effects, we also remove local search and reassess each component: as shown in Table 8, the replay buffer’s impact grows without search (11–19% drop vs. 4–6%), and eliminating augmented rewards still causes the largest performance degradation, underscoring their effectiveness in FoR.
Figure 3: Additional analysis on BlocksWorld. Left: Accuracy of FoR across different step settings with varying intermediate reward weight ($\lambda$). Right: Comparison of accuracy and diversity between SFT trained with varying data sizes and FoR trained on the test dataset.
Data-Efficiency. We test the final 20 examples from the 6-step test set and adjust the number of training examples for both FoR and SFT. As illustrated in the right plot of Figure 3, SFT’s accuracy improves with additional training data, while its diversity remains stable. However, SFT’s performance remains lower than FoR’s under any amount of training data. This is attributed to FoR’s ability to learn from divergent reasoning trajectories, enhancing trajectory coverage and improving generalization to new cases.
Discussion. FoR achieves higher reasoning accuracy and diversity by leveraging its exploratory nature (Madan et al., 2024) to sample multiple valid trajectories, covering a broader solution space without getting stuck on incorrect paths like SFT. This increases the likelihood of reaching a correct solution. A detailed case study that compares FoR with SFT is provided in Appendix G. Furthermore, FoR demonstrates robust performance across varying training set sizes, different LLM families, and larger model scales (see Appendix D).
@@ 5 Conclusion
We introduce FoR, which efficiently trains LLMs to generate diverse, high-quality reasoning paths proportional to unnormalized rewards. By leveraging a flow-based multi-step reasoning formulation, FoR adapts and improves GFlowNets training strategies. Across six reasoning tasks, FoR outperforms finetuning- and prompting-based baselines in both accuracy and diversity. Limitations are discussed in Appendix I.
Impact Statement
This study introduces FoR, a framework that enhances LLMs’ ability to solve multi-step reasoning problems. This approach lays the foundation for scalable, principled LLM reasoning, with potential applications in automated theorem proving, robotics, and scientific discovery, contributing to more robust and interpretable AI systems.
@@ Appendix A Additional Related Work
Reasoning with LLM. Recent LLMs (Achiam et al., 2023; Touvron et al., 2023; Bai et al., 2022; Chowdhery et al., 2023) have demonstrated great potentials in tackling complex reasoning tasks (Cobbe et al., 2021; Mishra et al., 2022; Hendrycks et al., 2021; Rein et al., 2023; Mialon et al., 2023). (1) Fine-tuning LLMs is a primary way to enhance their reasoning abilities, including SFT and reward-maximization reinforcement learning (RL) approaches. SFT with large-scale and high-quality datasets of reasoning chains has proven very effective (Yu et al., 2023c; Yue et al., 2023; Yuan et al., 2024a). Various methods for constructing training samples have been proposed when ground truth reasoning chains are not available. For example, STaR (Zelikman et al., 2022) uses online sampling with self-correction to find positive samples. ReSTEM (Singh et al., 2023) and V-STaR(Hosseini et al., 2024) filter samples with external verifiers. RL techniques, particularly reward-maximizing policy optimization methods like PPO, are widely employed in LLMs (Ouyang et al., 2022; Bai et al., 2022; Havrilla et al., 2024; Luong et al., 2024). However, both maximum likelihood training (i.e. SFT) and reward-maximizing policy optimization (e.g., PPO) do not encourage models to generate diverse solutions. (2) prompting-based reasoning algorithms aim to better elicit the knowledge inside LLMs without tuning their parameters. Techniques such as CoT (Wei et al., 2022) and its variants (Chen et al., 2022; Li et al., 2024b; Zhang et al., 2023e; Zhou et al., 2022; Kojima et al., 2022) have improved LLM performance by enabling them to generate intermediate steps before arriving at a final answer. To provide reasoning more guidance , self-evaluation (Xie et al., 2024b; Shinn et al., 2024; Madaan et al., 2024) and reward models are introduced to enhance reasoning process (Uesato et al., 2022; Lightman et al., 2023) Besides, a more relevant series of works combine LLM reasoning capabilities with planning and search algorithms such as MCTS (Hao et al., 2023; Feng et al., 2023; Zhao et al., 2024), tree and graph search (Jung et al., 2022; Zhu et al., 2022; Yao et al., 2024; Besta et al., 2024; Yao et al., 2023). Moreover, recent studies turn to amortizing computation for reasoning paths (Yuan et al., 2023; Wu et al., 2024; Bansal et al., 2024), such as self-correct (Kumar et al., 2024; Saunders et al., 2022), self-improvement (Tian et al., 2024; Yuan et al., 2024b).
@@
GFlowNets. GFlowNets (Bengio et al., 2021) were originally proposed to learn policies for sampling from unnormalized distributions, with a primary motivation from scientific discovery (Jain et al., 2023a), which requires generating diverse high-reward samples (Shen et al., 2023b; Roy et al., 2023; Zhang et al., 2023b; Ma et al., 2024; Pan et al., 2023b), such as molecular generation (Koziarski et al., 2024; Kim et al., 2024a; Lu et al., 2024) and biological sequence generation (Ghari et al., 2023; Jain et al., 2022). Beyond the science domain, GFlowNets have also been applied in various downstream applications such as recommendation systems (Liu et al., 2023b), domain adaptation (Zhu et al., 2023), combinatorial optimization (Zhang et al., 2023a; Kim et al., 2024b) and explainability of deep neural networks (Li et al., 2023b). Additionally, GFlowNets have proven to be suitable for sampling from posterior distributions (Hu et al., 2023b; Deleu et al., 2022, 2024; Zhang et al., 2022). As a reinforcement learning method, prior works have incorporated intermediate feedback with GFlowNets to address sparse reward issues (Pan et al., 2023a; Jang et al., 2023; Pan et al., 2022) and multi-objective rewards (Jain et al., 2023b; Hernandez-Garcia et al., 2023; Chen & Mauch, 2023). There are also theoretical analyses treating GFlowNets as recurrent MCMC (Deleu & Bengio, 2023) and variational inference (Malkin et al., 2022b; Zimmermann et al., 2022) that are used to model the distribution over trajectories.
@@
Diverse Thinking. Vertical and lateral thinking (Waks, 1997a; Ismayilzada et al., 2024) are two distinct approaches that differ significantly in their focus and methodology. Vertical thinking emphasizes logical, structured, and sequential reasoning, often following a step-by-step approach to solve problems. Our work aligns with this paradigm to generate multiple correct, structured reasoning processes to achieve specific goals. In contrast, lateral thinking prioritizes creativity and innovation, encouraging the exploration of unconventional perspectives and challenging established assumptions. There are multiple works that evaluate (Huang et al., 2023; Chen et al., 2024a; Kraaijveld et al., 2024; Todd et al., 2024) and promote(Zhong et al., 2024; Summers-Stay et al., 2023; Jiang et al., 2023; Lin et al., 2021) the lateral thinking ability of LLMs.
@@ Appendix B Preliminaries and Background
GFlowNets (Bengio et al., 2021, 2023; Liu et al., 2023a) are a class of models that amortize the cost of sampling from an intractable target distribution over terminal states $\mathcal{X}$ by learning a neural network-facilitated approximation of the target distribution using its unnormalized density or reward function. The task of sampling from this distribution resorts to a decision-making process. Below, we introduce GFlowNets with more details.
Settings. We are given a pointed directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$, where $S$ is a finite set of vertices (states), and $\mathcal{A}\subseteq\mathcal{S}\times\mathcal{S}$ is a set of directed edges (actions). If $s\to s^{\prime}$ is an action, we say $s$ is a parent of $s^{\prime}$ and $s^{\prime}$ is a child of $s$. There is exactly one state that has no incoming edge, called the initial state $s_{0}\in S$. States that have no outgoing edges are called terminal states. We denote by $\mathcal{X}$ the set of terminal states. A complete trajectory is a sequence $\tau=(s_{0}\to\ldots\to s_{n})$ such that each $s_{i}\to s_{i+1}$ is an action and $s_{n}=x\in\mathcal{X}$. We denote by $\mathcal{T}$ the set of complete trajectories and the terminal state as $\tau_{x}$.
@@
Training. Training GFlowNets considers achieving a consistent flow (Bengio et al., 2023; Malkin et al., 2022a), which means the flow for the forward direction should equal to the flow for the backward direction. Below we introduce relevant objectives.
Detailed Balance (DB). The DB objective (Bengio et al., 2023) requires learning two objectives in addition to parametric forward policy $P_{F}(\cdot|s)$: 1. A Backward policy, which is distribution $P_{B}(s^{\prime}|s;\theta)$ over the parents of any non-initial state. 2. A State flow function: $F(\cdot;\theta):\mathcal{S}\to\mathbb{R}_{>0}$. Then DB loss for a single transition $s\to s^{\prime}$ is defined as:
@@
Trajectory Balance (TB). Trajectory balance (Malkin et al., 2022a) introduces a backward policy $P_{B}$, which is a learned distribution $P_{B}(\cdot|s^{\prime})$ over the parents of every state $s\in S$. We have the following backward probability:
$P_{B}(\tau)$ $=P_{B}(s_{0}\to\ldots\to s_{n}|s_{n})=\prod_{i=0}^{n-1}P_{B}(s_{i}|s_{i+1})\quad\forall\tau\in\mathcal{T}\quad$ (11)
Then we give the trajectory balance constraint for any complete trajectory $\tau=(s_{0}\to\ldots\to s_{n})$:
$F(s_{0})\prod_{i=1}^{n}P_{F}(s_{t}|s_{t-1})=F(s_{n})\prod_{i=1}^{n}P_{B}(s_{i-1}|s_{i})$ (12)
where we have $P(s_{n})=F(s_{n})/Z$. To enforce this constraint, we convert equation 12 into an objective for optimization. Suppose that a model with parameters $\theta$ outputs estimated forward policy $P_{F}(\cdot|s;\theta)$ and backward policy $P_{B}(\cdot|s;\theta)$ for states $s$, as well as a scalar $Z_{\theta}$ estimating $F(s_{0})$. GFlowNets aim to approximate a Markovian Flow $F_{\theta}$ that $F(s_{n})=R(s_{n})$. For a complete trajectory $\tau=(s_{0}\to\ldots\to s_{n})$, define the trajectory balance loss as follows:
@@
If $\mathcal{L}_{TB}$ is made equal to 0 for every complete trajectory $\tau$, then equation 9 satisfies for all $x\in X$ and $Z$ is the inverse constant of proportionality: $Z=\sum_{x\in\mathcal{X}}R(x)$.
Conditional GFlowNets. In a GFlowNet, both the policy and reward function can be conditioned on additional information. For instance, in the tasks we focus on, a GFlowNet policy generates actions sequentially for an embodied reasoning problem, starting from an initial state $s_{0}$ and a goal $g$. Furthermore, the allowable actions vary depending on the specific $s_{0}$ in each case. The conditional GFlowNets we train achieve amortization by sharing the policy model across different $s_{0}$ and $g$, enabling the model to generalize to initial states and targets that were not seen during training.
@@ C.3 Efficiency Analysis
FoR Setup. During the training, we finetune the LLM with LoRA (Hu et al., 2021) with $r=32$, $\alpha=64$, and dropout=0.1. We set $\epsilon$ from 0.3 and decrease it to 0.01, $\beta$ from 1 to 2, and the probability $\delta$ using replay buffer increases from 0.3 to 0.5 throughout the iterations linearly. The learning rate is set to 1e-4 with a cosine annealing schedule, and the number of training iterations is set to 10. Reward weight $\lambda$ is set to 1.5. In our ablation study when setting $\lambda=0$, we add a small number $b=0.5$ to avoid $\log 0$. Table 7 shows the template we use for the forward policy in the 6-step setting, and its difference between 2-step and 4-step is only replacing the 6-step demonstration to 2-step and 4-step. During testing, we sample 8, 20, and 40 trajectories for 2, 4, and 6 steps respectively. As long as one trajectory reaches the goal, we label this instance as solved, all the baselines conform to the same rule.
Additional details for baselines. We compare FoR the following baselines:
@@
(5) Proximal Policy Optimization (PPO) (Schulman et al., 2017): This is a widely-used reward-maximization reinforcement learning method for LLM training. We design the objective to encourage the LLM to generate solutions that satisfy the goal. Following the common practice of previous work (Ouyang et al., 2022; Wang et al., 2023a), we penalize the policy if it deviates too much from the reference policy. Formally, the objective is $\max_{\pi_{\theta}}\mathbb{E}_{\tau\sim\pi_{\theta}}\left[R(x,y)\right]-\beta\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}(y\mid x)\|\pi_{\text{ref }}(y\mid x)\right]$.
@@
CoT (1-shot) 9.52 1.00 3.12 2.02 1.00 0.00
CoT (5-shot) 14.28 1.00 3.12 12.12 1.08 3.45
CoT (15-shot) 11.90 1.00 3.12 8.08 1.00 0.00
@@
Table 11: Baseline results with diversity-encouraging instruction prompt. "+" denotes the performance improvement compared to the absence of the prompt.
@@
Performance on OOD settings. We further assess performance on out-of-distribution (OOD) settings. Specifically, we train the model using FoR and SFT on a 2-step training set and evaluate it on a 4-step test set, and then train the model on the 4-step training set and evaluate it on the 6-step test set. This allows us to analyze their generalization on OOD problems. For prompting-based baselines, we use 2-step and 4-step examples as demonstrations, respectively.
@@
Additional results on search-based approaches. To assess whether search-based methods can generate diverse solutions, we execute ToT (DFS) multiple times (equal to the number of runs performed for our method) and compare their performance. The results shown in Table 13 demonstrate that though search-based methods improve the diversity, they still underperform FoR.
Table 13: Performance of Search-based methods with multiple runs. ${}^{\mathord{\clubsuit}}$ denotes results from a single run.
Method 2-step 4-step 6-step
Acc. (%) Acc. (%) Diversity Acc. (%) Diversity
ToT (${}^{\mathord{\clubsuit}}$ 13.33 16.67 - 8.08 -
ToT (DFS) 40.00 42.85 1.0 31.31 1.06
FoR 100.00 98.41 1.27 78.44 1.33
Additional baseline results with diversity-encouraging instruction. To further stimulate the diverse reasoning ability in the baseline approaches, we add a diversity-encouraging prompt as instruction:
@@
Additional case study. In Table 12, we show examples generated by FoR. We observe that after training, FoR can sample the terminal state with probability approximately proportional to the rewards, leading to an approximate sampling of different plans with the same probability. This empirically verifies the efficacy of the training objective.
@@
FoR Setup. See Figure 8 for the prompt template used in the experiment of the Game of 24. We use LoRA to train the model with $r=8$, $\alpha=32$, dropout=0.1. We load the LLM in fp16, and set the hyperparameters as follows: batch size = 4, learning rate = 1e-5, number of epochs = 5, and the reward weight $w=100$.